OpenStudy (anonymous):

f(x)=(x+3)^2(x-1)^2 Need to find x and y intercept power function of f max # turning points behavior for x intercept 1 and -3

sam (.sam.):

to find x and y intercept, set x=0 and y=0, When x=0, y=(0+3)^2(0-1)^2 y=9 when y=0. x=1

sam (.sam.):

To find power function of f, expand x^(2)-2x+1

sam (.sam.):

Since its a quadratic, the graph has 1 turning point

OpenStudy (anonymous):

So my pwer of f would be 1 right?

sam (.sam.):

power function is 2, because x^(2)-2x+1 Observe the term with the greatest degree. x^2 Now, you know the power function is f(x) = x^2

OpenStudy (anonymous):

How do I fiqure out the behavior of the graph near x-intercept 1 and x-intercept -3?

sam (.sam.):

x-intercept 1 and x-intercept -3 , means the curve touches the x axis 2 times,